OpenStudy (anonymous):

Find the type and number of solutions for the equation. 2x – x2 = 4

OpenStudy (anonymous):

convert to this form x2 - 2x + 4 = 0 then find the Discriminant (b2 - 4ac) = 4 - 16 = -8 based on that the roots will be complex numbers